# Interpreting token compositionality in LLMs: A robustness analysis

### v2

@@ Abstract
Nura Aljaafari1,4†,  Danilo S. Carvalho1,3,  André Freitas1,2,3 1 Department of Computer Science, University of Manchester, United Kingdom 2 Idiap Research Institute, Switzerland 3 National Biomarker Centre, CRUK-MI, Univ. of Manchester, United Kingdom 4 King Faisal University, Al Hofuf, KSA {firstname.lastname}@[postgrad.]manchester.ac.uk
@@ 1 Introduction
The number and diversity of applications for Transformer-based large language models (LLMs) are currently at an accentuated growth, considering their performance at major NLP tasks continues to increase. Yet, the understanding of characteristics linked to LLMs critical deficiencies, such as hallucinations and lack of interpretability, remain limited. In particular, there is the matter of linguistic compositionality: how different units of text (morphemes, words, phrases) are combined into units of meaning, and how this relates to models of language interpretation.
Figure 1: Illustration of the CAP process. Constituent Segmentation creates segments based on token constituents, representing linguistic units like words or phrases. CAP is applied at layer $m$, pooling activations using aggregation protocols (e.g., max, mean, sum). CAP can be applied to any component, provided dimensionality is maintained. Activations at layer $m+1$ show the model’s new-reduced dimensionality. The results graph reflects how CAP is applied at different depths, with accuracy recorded for comparison.
Uncovering compositional representations in Transformer-based models is a key step in reducing the semantic gap between user intent and model interpretation, explaining model behaviour in different settings and mitigating hallucinations and inconsistent reasoning. To this effect, current work has predominantly focused on mapping input-output pairs Yin et al. (2024), input embeddings Haslett (2024) and layer-wise outputs Yu and Ettinger (2020); Modarressi et al. (2023) to expected latent representations by means of semantic similarity or correlation. Such focus reflects two intuitive expectations regarding LLM behaviour:
  • LLMs should have an internal representation that surfaces a token/word-level compositionality.
  • While substantially distributed across the model’s parameters, (phrasal, sentence-level) meaning is expected to be “localisable” in order to be addressed during inference.
Prior work has shown, through different perturbation experiments, that LLM representations and responses are fragile and unreliable across models Wang et al. (2023); Fodor et al. (2024); Hu et al. (2024). Furthermore, analysis of phrasal representations Yu and Ettinger (2020); Carvalho et al. (2025) revealed no significant correlation with phrasal compositional semantics. However, the mechanisms behind such obfuscation of the knowledge addressable by a Transformer model are still open research questions, which motivates this work.
Aiming to elucidate how token representations, particularly with respect to compositionality, affect LLM interpretation, this work investigates the sensitivity/robustness of LLMs across local compositional perturbations at different stages of a Transformer model. In this study, we specifically focus on autoregressive decoder Transformers, investigating their compositional properties. We propose Constituent-Aware Pooling (CAP), a perturbation method for pooling LLM activations that correspond to individual tokens into activations representing more coherent linguistic units, using it to improve the understanding of compositionality mechanisms within Transformer-based LMs (for an overview of compositionality and localisation, see Appendix A).
Through this intervention, contrary to expectations of incremental compositional/semantic build-up, we found that LLMs focus heavily on isolated token features rather than hierarchically integrating semantic information across layers, leading to significant performance drops when CAP is applied. These fluctuations, especially pronounced in early and middle layers, indicate that aggregation of syntactic and semantic information is distributed across multiple layers rather than localised in any single layer. Surprisingly, larger models are more fragile to compositional perturbations than smaller ones, highlighting the significant difference in the way Transformer-based LMs build meaning representations.
Additionally, we interpret these empirical findings under the light of an information theoretical framework by suggesting that Transformer models maximise information gain about the next predicted token by delaying the aggregation of input token representations to later layers, thereby reducing mutual information (redundancy) between tokens at the same layer to optimise prediction, leading to longer aggregation paths across multiple layers.
In summary, the contributions of this work are:
1. A systematic analysis of compositional aggregation (word and phrasal-level) robustness for current LLMs.
2. A theoretical account for the observed results, which also explains the difficulty in locating compositional semantic representations in Transformer models.
Based on the empirical results and theoretical model presented here, we postulate that traditional compositional semantic representations cannot be isolated to any particular (intermediate) stage of a standard Transformer model. This is independent of model size, supervision type or inference task, but linked to the number of hidden layers. Our findings point toward the use of specialised architectures and/or training objectives in order to elicit such representations. The supporting datasets and software are available at a public repository111¡anonymised url¿.
@@ 3 Assessing compositional aggregation robustness
To accurately assess the effects of compositional grouping at different layers of abstraction within transformer models, the inference objective should be a task that is both: 1) strictly dependent on the input tokens and their composition, with few possible input variations; 2) contain as few tokens as possible in the output. For this reason, the following tasks were selected (Figure 1):
@@
This process consolidates activations for each syntactic unit, enabling systematic evaluation of compositional robustness across layers. For simplicity, we demonstrate the operation over these components, but this approach can be extended to any transformer’s components, provided that the dimensional requirements for information flow, as described in Elhage et al. (2021), are respected. For example, consider attention layer internal activations of shape $X\in\mathbb{R}^{B\times H_{a}\times K\times K}$, where $H_{a}$ is the number of attention heads, and $K$ represents the query and key token dimensions. Applying CAP with the Sum protocol involves aggregating activations over the query range $[s_{i},e_{i}]$ and the key range $[s_{j},e_{j}]$. The grouped activations are computed as: $\alpha([s_{i},e_{i}],[s_{j},e_{j}])=\sum_{t=s_{i}}^{e_{i}}\sum_{t^{\prime}=s_{j}}^{e_{j}}X[b,h,t,t^{\prime}]$. After applying CAP, the grouped activations have the shape $X\in\mathbb{R}^{B\times H_{a}\times G\times G}$, where $G$ is the number of grouped syntactic units. This ensures that query-key interactions are consolidated into cohesive syntactic units, aligning activations with higher-level linguistic structures. We examine CAP’s reduction ratio ($K\to G$) at the word-level and its effects across models, with detailed analysis in Appendix C.
@@ 4.1 Experimental setup & datasets
The methodology was tested across various decoder-only transformer models Vaswani (2017). Our main focus was on GPT-2 (small: 124M, medium: 355M, large: 774M parameters) Radford et al. (2019), Gemma (2B parameters) Team et al. (2024), Llama (3B, 8B, and 8B-instruct parameters) Dubey et al. (2024), and Qwen (0.5B, 1.5B, and 3B parameters) Yang et al. (2024). These models use different tokenisation approaches: byte-level BPE (GPT-2, Qwen), expanded BPE with 128K vocabulary (Llama3), and SentencePiece (Gemma). Models were tested before and after task-specific fine-tuning (3 epochs, learning rate 5e-5). This selection spans diverse architectures, sizes, and tokenisation strategies (see Appendix B.3 for further details on the models and fine-tuning parameters).
@@ 4.2 Results and discussion
Compositional inference in LLMs is not a purely incremental process. Contrary to expectations of a smooth and steady layer-wise performance improvement, we observe significant fluctuations when CAP is applied across layers (Figure 2). Performance drops notably in early and middle layers, followed by sharp improvements (Figure 2 (a)-(c), (e), and (f)), suggesting these layers struggle to process CAPed activations, particularly the pooled linguistic features captured in earlier layers. Rather than progressively building semantic information from individual tokens to complex phrases, the models appear to focus heavily on isolated token features. The results indicate that attention is distributed over input tokens and model layers in a non-systematic and decentralised manner that is highly context-dependent, showing minimal reliance on sequential or positional relationships of constituents. This phenomenon is particularly evident in the sharp decline in SP and HP tasks, where contextual information is limited during phrase-level CAP application. We argue that this behaviour stems from the model’s training objective, which maximises information gain in each layer towards predicted tokens at the cost of reducing mutual information between tokens in a single layer. This behaviour means that aggregation, including syntactic, is performed across multiple layers and thus is not localisable from any single given layer. An information theoretical analysis elaborates this reasoning in Section 5. Our findings highlight how compositional structures are highly sensitive to token representation dynamics across layers, suggesting that performance fluctuations can be attributed to information loss incurred as a function of token mutual information across layers.
@@
Larger models are more fragile to compositional perturbations. The IDM task highlights this fragility in larger models, as larger models rely on finer feature extraction. Within families, distinct patterns emerge: original Qwen’s smaller variants show better IDM robustness (e.g., at position 25% there was a 7.69% drop on Qwen-1.5B vs 12.11% on Qwen-3B), while Llama3 exhibits capacity-dependent behaviour with the 3B variant being more vulnerable than 8B. Notably, Llama3-8B-Instruct, despite being fine-tuned, performed worse in IDM compared to Llama 8B and 3B, but excels in structured tasks (SP, HP), underscoring the role of architecture and training in compositional robustness. Despite having similar reduction ratios to Llama models (see Appendix C), Gemma-2B shows greater sensitivity to perturbations (e.g., at position 1% Max: Gemma-2B drops 97.91% vs. Llama3-8B’s 25.49%), likely due to its larger vocabulary enabling finer-grained tokenisation. While fine-grained token knowledge benefits standard tasks, it appears to increase susceptibility to compositional perturbations. The superior performance of Llama3-8B over its 3B variant can be attributed to its enhanced capacity for maintaining feature relationships across layers while preserving key compositional information. While larger models excel in standard tasks (see Appendix E.1), they exhibit a greater reliance on the identification of intrinsic features in the early layers. We find that phrasal-level CAP substantially impacts Gemma-2B and Llama models, suggesting a heavy dependence on layer-wise information gain, where they separate features in an uncorrelated and highly distinct manner. While this aids in identifying complex feature patterns, it also makes them more vulnerable to contextual noise—a weakness that threatens their robustness and integrity. Notably, Qwen models outperform Llama and Gemma despite similar parameter counts, likely due to byte-level BPE tokenisation and multilingual training, which enhance compositional stability, whereas Llama’s expanded BPE and Gemma’s SentencePiece prioritise efficiency over phrase retention, increasing vulnerability to CAP interventions.
@@ 5 Information Gain & Token Mutual Information
where $p$, $q$ and $r$ are layer indices, i.e., subsequent layers have more information about the inputs than previous ones. This can be explained in that optimising Equation 5 can be achieved by retaining at each $R_{l_{p}}(X)$ only the necessary information to maximise $\sum_{i,j}{IG_{R_{l_{q}}(X_{i}),MHA(R_{l_{p}}(X_{j}))}}$, where $MHA(R_{l_{p}}(X_{j}))$ is the multi-head attention weighted representation. Such an objective implies minimising the mutual information $I(R_{l_{p}}(X_{i}),R_{l_{p}}(X_{j}))$, i.e., reducing redundancy across tokens from the same layer. Therefore, token dependencies will tend to be modelled by aggregation paths spanning multiple layers, with more layers allowing for more complex and longer paths. This is in line with the findings of Mechanistic Interpretability studies Elhage et al. (2021); Conmy et al. (2023). Equation 6 also implies that the earlier an aggregation is done, the larger the impact it will have on $IG_{Y,X}$, which explains the empirical results.
The effects of $I(R_{l}(X_{i}),R_{l}(X_{j}))$ on LLMs are further compounded by the tokenisation objective (e.g., BPE, WordPiece), which minimises $I(X_{i},X_{j})$, i.e., token redundancy, as a means of reducing the vocabulary size, leading to longer aggregation paths.
@@ 6 Related work
General compositionality. Compositionality, a fundamental concept across disciplines, concerns understanding complex systems through their simpler components’ meanings. It spans philosophy, cognitive science, linguistics, and machine learning Tull et al. (2024). In ML and AI, compositionality enables models to generalise, reason about, and interpret complex meanings from basic components. In cognitive science, the Language of Thought Hypothesis (LOTH) Fodor (1975) proposes that mental processes are compositional Rescorla (2024). Recent ML studies explore how transformers and neural networks build inferences. The logit lens Nostalgebraist (2020) demonstrated that transformers build predictions progressively where early layers make initial guesses and deeper layers refine guesses with broader context. Dai et al. (2022) show feed-forward layers act as key-value memories, combining information for complex predictions. MEMIT Meng et al. (2023) and PMET Li et al. (2025) show how controlled inferences can be built by manipulating models’ components. These studies suggest transformers may employ compositional processing, systematically building complex representations from simpler components.
Linguistic compositionality. In linguistics, compositionality is often linked to the principle of compositionality, which states that the meaning of a complex expression is derived from "the meanings of its constituent parts, the grammar used to combine them, and the syntactic structure as a whole" Montague and Thomason (1975). Yu and Ettinger (2020) found that transformers mainly encode individual word content rather than true phrase-level meaning, with model performance often tied to word overlap rather than compositional understanding. Carvalho et al. (2025) probed transformer models’ representation of adjectival modifier phenomena in adjective-noun phrases. Their tests revealed that models partially capture meaning intersective composition, which does not generalise across other adjective types. DecompX Modarressi et al. (2023) traces individual token representations through transformer layers, enabling analysis of compositional behaviour at the level of specific token predictions without requiring cross-layer vector aggregation. Haslett (2024) demonstrate that models from Multilingual BERT through GPT-4 frequently fail to segment words meaningfully, especially in non-Latin scripts where morpheme boundaries are poorly captured.
This work complements and extends these findings to characterise the fundamental disconnect between transformer representations and compositional semantics observed across aforementioned studies.

### v3

@@ Abstract
Nura Aljaafari1†,  Danilo S. Carvalho1,3,  André Freitas1,2,3 1 Department of Computer Science, University of Manchester, United Kingdom 2 Idiap Research Institute, Switzerland 3 National Biomarker Centre, CRUK-MI, Univ. of Manchester, United Kingdom {firstname.lastname}@[postgrad.]manchester.ac.uk
@@ 1 Introduction
Large language models (LLMs) based on Transformer architectures have rapidly expanded in scope and capability, demonstrating strong performance across a wide range of NLP tasks. However, critical limitations remain, including hallucinations, poor interpretability, and limited semantic transparency. One open challenge concerns linguistic compositionality: how models combine smaller units of text (e.g., morphemes, words, phrases) into coherent meaning structures, and how this process is reflected in internal representations.
Figure 1: Illustration of the CAP process. Constituent segmentation identifies linguistic units (e.g., words or phrases), and CAP pools their activations at layer $m$ using aggregation (e.g., max, mean, sum). This operation reduces sequence length, and the modified activations are propagated to layer $m{+}1$. The results graph shows task accuracy under CAP at different depths.
Understanding how and where compositional structure is encoded in LLMs is essential for bridging the gap between user intent and model behaviour. Prior work has explored this by aligning model inputs and outputs Yin et al. (2024), embedding spaces Haslett (2024), or layer-wise activations Yu and Ettinger (2020); Modarressi et al. (2023) with expected semantic representations. These approaches are grounded in two intuitive assumptions: (1) that LLMs internally represent compositional structure at the token or word level, and (2) that this information should be at least partially localisable at specific layers during inference.
Several studies have revealed that LLMs are often brittle under perturbation Wang et al. (2023); Fodor et al. (2024); Hu et al. (2024), and that phrase-level representations may fail to align with expected semantics Carvalho et al. (2025). Despite this, the mechanisms behind such fragility, particularly at the level of internal activations, remain poorly understood.
To investigate this, we propose Constituent-Aware Pooling (CAP), a structured perturbation method that groups token-level activations into larger constituent units (e.g., words or phrases) at arbitrary layers. CAP enables systematic probing of whether, and where, semantic meaning is robustly composed within the model. By applying CAP at varying depths, we assess the fragility of internal representations to compositional perturbations and examine whether, and how, semantic abstraction is distributed across layers.
Our empirical findings challenge common assumptions of hierarchical semantic buildup. Rather than gradually constructing compositional meaning across layers, LLMs often retain token-level focus well into the middle layers. Applying CAP, even at semantically coherent groupings, results in substantial accuracy degradation, especially in earlier layers. Surprisingly, larger models are more sensitive to such perturbations than smaller ones, suggesting increased representational fragility with scale.
We contextualise these results using an information-theoretic lens, proposing that Transformer models delay aggregation to maximise token-level information throughput. This leads to distributed, rather than localised, composition across layers, resulting in longer dependency paths and reduced mutual redundancy at each layer.
In summary, our contributions are:
  • A systematic analysis of how current LLMs handle constituent-level composition, evaluated via CAP across layers, models, and tasks.
  • A theoretical explanation grounded in information theory, suggesting that LLMs optimise for prediction by postponing semantic integration, thus fragmenting compositional meaning across depth.
We conclude that compositional semantics are not reliably localisable within any fixed layer of standard Transformer models. This holds across model scales, supervision types, and inference tasks, and instead appears tied to architectural depth. Our results suggest that recovering explicit compositional structure may require specialised training objectives or architectural constraints. Supporting code and datasets are available at a public repository111¡ anonymised url¿.
@@ 3 Assessing compositional aggregation robustness
To accurately assess the effects of compositional grouping at different layers of abstraction within transformer models, the inference objective should be a task that is both: 1) strictly dependent on the input tokens and their composition, with few possible input variations; 2) contains as few tokens as possible in the output. For this reason, the following tasks were selected (Figure 1):
@@
This process consolidates activations for each syntactic unit, enabling systematic evaluation of compositional robustness across layers. For simplicity, we demonstrate the operation over these components, but this approach can be extended to any transformer’s components, provided that the dimensional requirements for information flow, as described in Elhage et al. (2021), are respected. For example, consider attention layer internal activations of shape $X\in\mathbb{R}^{B\times H_{a}\times K\times K}$, where $H_{a}$ is the number of attention heads, and $K$ represents the query and key token dimensions. Applying CAP with the Sum protocol involves aggregating activations over the query range $[s_{i},e_{i}]$ and the key range $[s_{j},e_{j}]$. The grouped activations are computed as: $\alpha([s_{i},e_{i}],[s_{j},e_{j}])=\sum_{t=s_{i}}^{e_{i}}\sum_{t^{\prime}=s_{j}}^{e_{j}}X[b,h,t,t^{\prime}]$. After applying CAP, the grouped activations have the shape $X\in\mathbb{R}^{B\times H_{a}\times G\times G}$, where $G$ is the number of grouped syntactic units. This ensures that query-key interactions are consolidated into cohesive syntactic units, aligning activations with higher-level linguistic structures. We examine CAP’s reduction ratio ($K\to G$) at the word-level and its effects across models, with detailed analysis in Appendix C. We refer the reader to Appendix B.4 for further details on how CAP affects sequence length and interacts with positional encodings.
@@ 4.1 Experimental setup & datasets
The methodology was tested across various decoder-only transformer models Vaswani (2017). Our main focus was on GPT-2 (small: 124M, medium: 355M, large: 774M parameters) Radford et al. (2019), Gemma1 (2B parameters) Team et al. (2024), Llama (3B, and 8B parameters) Dubey et al. (2024), and Qwen (0.5B, 1.5B, and 3B parameters) Yang et al. (2024). These models use different tokenisation approaches: byte-level BPE (GPT-2, Qwen), expanded BPE with 128K vocabulary (Llama3), and SentencePiece (Gemma). Models were tested before and after task-specific fine-tuning (3 epochs, learning rate 5e-5). This selection spans diverse architectures, sizes, and tokenisation strategies (see Appendix B.3 for further details on the models and fine-tuning parameters).
@@ 4.2 Results and discussion
Compositional inference in LLMs is not a purely incremental process. Contrary to expectations of a smooth and steady layer-wise performance improvement, we observe significant fluctuations when CAP is applied across layers. Performance drops notably in early and middle layers, followed by sharp improvements (Figure 2 (a)-(c), (e), and (f)), suggesting these layers struggle to process CAPed activations, particularly the pooled linguistic features captured in earlier layers. Rather than progressively building semantic information from individual tokens to complex phrases, the models appear to focus heavily on isolated token features.
An important distinction arises between TW-CAP, which groups tokens according to model-specific tokenisation, and TP-CAP, which applies externally parsed syntactic structures. While TP-CAP introduces richer constituent information, it may not align with the model’s internal segmentation or syntactic reasoning. This misalignment is not a flaw in CAP, but rather a diagnostic signal: if LLMs encoded human-like syntax, TP-based grouping should be minimally disruptive. The observed drop in performance under TP-CAP suggests that LLMs do not consistently internalise hierarchical syntactic structures. This finding underscores the model’s emphasis on local token-level information and supports the conclusions drawn in our information-theoretic analysis.
The results indicate that attention is distributed over input tokens and model layers in a non-systematic and decentralised manner that is highly context-dependent, showing minimal reliance on sequential or positional relationships of constituents. This phenomenon is particularly evident in the sharp decline in SP and HP tasks, where contextual information is limited during phrase-level CAP application. We argue that this behaviour stems from the model’s training objective, which maximises information gain in each layer towards predicted tokens at the cost of reducing mutual information between tokens in a single layer. This behaviour means that aggregation, including syntactic, is performed across multiple layers and thus is not localisable from any single given layer. An information theoretical analysis elaborates this reasoning in Section 5. Our findings highlight how compositional structures are highly sensitive to token representation dynamics across layers, suggesting that performance fluctuations can be attributed to information loss incurred as a function of token mutual information across layers.
@@
Larger models are more fragile to compositional perturbations. The IDM task highlights this fragility in larger models, as larger models rely on finer feature extraction. Within families, distinct patterns emerge: original Qwen’s smaller variants show better IDM robustness (e.g., at position 25% there was a 7.69% drop on Qwen-1.5B vs 12.11% on Qwen-3B), while Llama3 exhibits capacity-dependent behaviour with the 3B variant being more vulnerable than 8B. Despite having similar reduction ratios to Llama models (see Appendix C), Gemma-2B shows greater sensitivity to perturbations (e.g., at position 1% Max: Gemma-2B drops 97.91% vs. Llama3-8B’s 25.49%), likely due to its larger vocabulary enabling finer-grained tokenisation. While fine-grained token knowledge benefits standard tasks, it appears to increase susceptibility to compositional perturbations. The superior performance of Llama3-8B over its 3B variant can be attributed to its enhanced capacity for maintaining feature relationships across layers while preserving key compositional information. While larger models excel in standard tasks (see Appendix E.1), they exhibit a greater reliance on the identification of intrinsic features in the early layers. We find that phrasal-level CAP substantially impacts Gemma-2B and Llama models, suggesting a heavy dependence on layer-wise information gain, where they separate features in an uncorrelated and highly distinct manner. While this aids in identifying complex feature patterns, it also makes them more vulnerable to contextual noise—a weakness that threatens their robustness and integrity. Notably, Qwen models outperform Llama and Gemma despite similar parameter counts, likely due to byte-level BPE tokenisation and multilingual training, which enhance compositional stability, whereas Llama’s expanded BPE and Gemma’s SentencePiece prioritise efficiency over phrase retention, increasing vulnerability to CAP interventions.
@@ 5 Information Gain & Token Mutual Information
where $p$, $q$ and $r$ are layer indices, i.e., subsequent layers have more information about the inputs than previous ones. This can be explained in that optimising Equation 5 can be achieved by retaining at each $R_{l_{p}}(X)$ only the necessary information to maximise $\sum_{i,j}{IG_{R_{l_{q}}(X_{i}),MHA(R_{l_{p}}(X_{j}))}}$, where $MHA(R_{l_{p}}(X_{j}))$ is the multi-head attention weighted representation. Such an objective implies minimising the mutual information $I(R_{l_{p}}(X_{i}),R_{l_{p}}(X_{j}))$, i.e., reducing redundancy across tokens from the same layer. Therefore, token dependencies will tend to be modelled by aggregation paths spanning multiple layers, with more layers allowing for more complex and longer paths. This is in line with the findings of Mechanistic Interpretability studies Elhage et al. (2021); Conmy et al. (2023). Equation 6 also implies that the earlier an aggregation is done, the larger the impact it will have on $IG_{Y,X}$, which explains the empirical results. The effects of $I(R_{l}(X_{i}),R_{l}(X_{j}))$ on LLMs are further compounded by the tokenisation objective (e.g., BPE, WordPiece), which minimises $I(X_{i},X_{j})$, i.e., token redundancy, as a means of reducing the vocabulary size, leading to longer aggregation paths.
@@ 6 Related work
Compositionality, the principle that the meaning of complex expressions is derived from their parts and structure, is foundational in linguistics, cognitive science, and AI Fodor (1975); Montague and Thomason (1975); Tull et al. (2024). In neural models, compositionality enables generalisation and interpretability, yet remains difficult to diagnose and enforce Donatelli and Koller (2023). Several studies investigate how and where compositional representations emerge in transformer models. Carvalho et al. (2025) observed similar effects in adjective-noun phrase probing, while Haslett (2024) found that models struggle to segment or represent morphemes, especially in non-Latin scripts, suggesting breakdowns in both form and meaning composition. The logit lens Nostalgebraist (2020) demonstrated that transformers build predictions progressively where early layers make initial guesses and deeper layers refine guesses with broader context. Dai et al. (2022) show feed-forward layers act as key-value memories, combining information for complex predictions. MEMIT Meng et al. (2023) and PMET Li et al. (2025) show how controlled inferences can be built by manipulating models’ components. Some nuance emerges in later-layer behaviours. DecompX Modarressi et al. (2023) traced token representations layer-by-layer and observed partial shifts toward integration. Yu and Ettinger (2020) tested model encoding and found that transformers mainly encode individual word content rather than true phrase-level meaning. While some models appear more compositional under certain conditions, general trends remain unclear. For example, Dankers et al. (2022) demonstrate that models can show unexpectedly high or low compositionality depending on the data and task, suggesting exposure and framing affect outcomes as much as architecture. Petty et al. (2024) show that deeper Transformers tend to generalise more compositionally than shallower ones, though the benefits diminish beyond a certain depth. This highlights that architectural depth, not just scale, may shape compositional ability, though with diminishing returns. In multi-step reasoning tasks, models often fall back on shallow pattern matching rather than true decomposition Dziri et al. (2023).
Prior work has primarily relied on synthetic tasks to assess compositional generalisation, focusing on properties such as systematicity, productivity, and substitutivity Hupkes et al. (2020); Lake and Baroni (2018), these setups often abstract away from the complexities of natural language. More recent studies using natural data are often limited to small domains such as semantic parsing or machine translation Lake and Baroni (2018); Kim and Linzen (2020), and typically lack insight into internal representations.
In contrast to prior works focused on final outputs or synthetic tasks, CAP is a method for probing compositional structure within LLMs using real inputs. It intervenes directly on hidden activations, merging token-level representations into word- or phrase-level constituents at various depths. This allows us to evaluate where semantic composition occurs and how robust LLMs to structured perturbations. Unlike surface-level probes, CAP provides a targeted, activation-level lens on how meaning is constructed and distributed across model layers and linguistic units.